Archive for May, 2007

Additional Capacity Live

Posted: Friday, May 25th, 2007 at 14:25 by Steve Lalonde

We have just been informed the BT have enabled the next 155 on central4 :)

Emergency Rebalance

Posted: Wednesday, May 23rd, 2007 at 17:45 by Steve Lalonde

I am going to rebalance the central pipes at 18:00 today.

We have 2 totally full and 1 almost empty and 1 just right, I have tried to move a few users from the full pipes but its not working. so its last resort time.

  1. Steve Lalonde says:

    Well that did not help to solve the imbalance we ended up after everyone reconnected, back with almost exactly the same load distribution as before. It did however prove that BT have a problem now acknowledged by a senior BT engineer. I’m now raising this as a fault with BT.

LTS Software upgrade

Posted: Thursday, May 17th, 2007 at 16:58 by Neil Watson

Following on from today’s problems with L2TP sessions, it is necessary to perform an upgrade to the IOS software of our LTS pool. The reloads are scheduled for 03.00hrs and 04.00hrs on 24th May 2007 and are expected to take no longer than 25 minutes each. Multiple reloads have been scheduled to reduce the Radius load on our customer’s servers. During this period L2TP customers will experience tunnel drops. We apologise for any inconvenience this may cause.

Note: These reloads will only affect L2TP users, not standard Entanet Partners and their end users.

L2TP stuck session issues

Posted: Thursday, May 17th, 2007 at 11:21 by Neil Watson

We are currently experiencing a very high number of L2TP stuck sessions that are constantly trying to connect to our service, they are, however, not passing radius requests and hence cannot connect. In order to remove these sessions it is necessary to disconnect all users, however this will be carried out in a controlled manner, with each tunnel cleared in turn. Through our tests, it has already been evidenced that reconnection to a different tunnel resolves the issue for the end user. This controlled method of disconnection also ensures that L2TP providers Radius servers should not become overloaded with requests.

This work will be carried out over the course of today, as, given the number of affected users, it cannot be performed out of hours. We apologise for the inconvenience that will cause.

Please note that this work only affects L2TP users, not standard Entanet Partners and their end users.

Entanet Voice: VoIP voicemail server

Posted: Thursday, May 17th, 2007 at 09:58 by Wayne Hawtin

We are currently experiencing problems with the voicemail server which means that this service is currently unavailable. Our engineers are working on this and hope to have the problem resolved in the next few hours. Further updates will be made as soon as available.

Telehouse Metro

Posted: Saturday, May 12th, 2007 at 11:45 by Steve Lalonde

We are currently unable to contact our router in Telehouse Metro, engineers are working on the problem.

  1. Steve Lalonde says:

    telehouse-metro.core is now back in service, we will be analyzing the reason for the crash with Cisco.

smtp2.enta.net

Posted: Thursday, May 10th, 2007 at 10:25 by Neil Watson

We are currently experiencing an issue with outbound e-mail delivery, with some customers using smtp.enta.net for outbound mail experiencing intermittent problems when attempting to send email. We currently have engineers tasked on this issue and the cause of the problem has been identified. Our systems engineers are working on a resolution with priority. Apologies for any inconvenience this may be causing you. Full service will be resumed shortly.

  1. Dale Lloyd says:

    Full service is now restored, apologies for any inconvenience this may have caused.

Central Pipe Rebalance

Posted: Wednesday, May 2nd, 2007 at 07:21 by Steve Lalonde

Due to the massive inbalance on all centrals I am going to reset all connections. Central 1 and 2 are completely full while 3 and 4 are almost empty. This affects all directly connected sessions but not L2TP sessions. Looks like the inbalance was caused by BT maintenance in the early hours of yesterday morning (still trying to confirm this).